OpenStudy (anonymous):

How do you multiply radicles (square roots)? For example: 2√5 times 2√5, √10 times √12, 3√8 times 4√13

OpenStudy (anonymous):

all you do is multiply them lol 3*8 4*13 like that

OpenStudy (anonymous):

Wait what so the answer is 24, 52... that makes no sense

OpenStudy (lgbasallote):

\[\large 2\sqrt 5 \times 2\sqrt 5 = ( 2\times 2) \times (\sqrt 5 \times \sqrt 5)\] \[\large =4 \times (\sqrt 5 \times \sqrt 5) = 4 \times 5 = 20\]

OpenStudy (anonymous):

:O Thank you. what if you have 2 different square roots such as 2√5 x 7√9. Is it (2*7)+(5*9)=14√45...